OpenStudy (anonymous):

in a regular polygon, the ratio of the measure of the exterior angle to the measure of its adjacent interior angle is 1 to 4. how many sides does the polygon have?

OpenStudy (nali):

can u plz draw a picture

OpenStudy (anonymous):

there is no picture thats what i have to do...

OpenStudy (nali):

The sum of the interior angles in degrees of any polygon is given by: 180(n-2) where n is the number of sides of the polygon. The sum of the exterior angles of any polygon is 180 degrees. So if you have a regular n-gon where one of the interior angles is 4 times larger than the adjacent exterior angle, then the sum of the interior angles must be 4 times the sum of the exterior angles, so: 180(n-2) = 4 (180) Just solve for n here is my source http://www.algebra.com/algebra/homework/Polygons/Polygons.faq.question.196757.html

OpenStudy (anonymous):

i thought the total sum of exterior angles is 360

OpenStudy (nali):

so do u get it though

OpenStudy (nali):

there are 6 sides in the polygon

OpenStudy (anonymous):

kinda...ill put it together at one point

OpenStudy (nali):

so do u need additional help

OpenStudy (anonymous):

i swaer the totoal sum of exterior angle is always 360, only a triangles sum is 180

OpenStudy (nali):

imagine that is is a hexagon  144 +36 = 180